About Me With Graphics

Marek Rychlik

University of Arizona, Department of Mathematics February 8, 2005

1 This section will have a fancy figure

Figure 1: This a caption

Figure 2: This a caption

Area(Circle) = πr^{2}

2

[figure: Area(Circle) = πr^{2}]

Figure 3: This is pure L^{A}TEX